Preble Company manufactures one product. Its variable manufacturing overhead is applied to production based on direct labor-hours and its standard cost card per unit is as follows:





Direct material: 6 pounds at $8.00 per pound$48.00
Direct labor: 3 hours at $14 per hour
42.00
Variable overhead: 3 hours at $5 per hour
15.00
Total standard variable cost per unit$105.00




The company also established the following cost formulas for its selling expenses:



Fixed Cost per Month
Variable Cost per Unit Sold



Advertising$250,000




Sales salaries and commissions$200,000

$17.00
Shipping expenses



$8.00









The planning budget for March was based on producing and selling 19,000 units. However, during March the company actually produced and sold 24,000 units and incurred the following costs:


  1. Purchased 160,000 pounds of raw materials at a cost of $7.20 per pound. All of this material was used in production.

  2. Direct-laborers worked 60,000 hours at a rate of $15.00 per hour.

  3. Total variable manufacturing overhead for the month was $336,600.

  4. Total advertising, sales salaries and commissions, and shipping expenses were $260,000, $480,000, and $165,000, respectively.

1. What raw materials cost would be included in the company’s flexible budget for March?

2. What is the materials quantity variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input the amount as a positive value.)

3. What is the materials price variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input the amount as a positive value.)

4. If Preble had purchased 175,000 pounds of materials at $7.20 per pound and used 160,000 pounds in production, what would be the materials quantity variance for March? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance.). Input the amount as a positive value.)
Answer #1

1. Raw materials cost in Flexible budget

= 24,000 units x 6 pounds x $8 per pound

=$1,152,000

 

 

2. Material quantity variance

= (Standard Quantity - Actual Quantity)* Standard price

= ( 144,000 -160,000) x $8

= 128,000 (U)

Where Standard Quantity = 24,000 units x 6 pound per unit =144,000 pounds

 

3. Material Price Variance

= ( Standard Price - Actual Price ) x Actual Quantity

= ($8 -$7.2) x 160,000  

=128,000 (F)

 

4.  Material quantity variance

= (Standard Quantity - Actual Quantity consumed)* Standard price

= ( 144,000 -160,000) x $8

= 128,000 (U)

Where Standard Quantity = 24,000 units x 6 pound per unit =144,000 pounds

 


answered by: sidjn50